OpenStudy (anonymous):

How do I do this???? The amount of money in an account with continuously compounded interest is given by the formula mc009-1.jpg, where P is the principal, r is the annual interest rate, and t is the time in years. Calculate to the nearest hundredth of a year how long it takes for an amount of money to double if interest is compounded continuously at 3.8%. Round to the nearest tenth.

OpenStudy (anonymous):

***Formula is A = Pe^rt

OpenStudy (anonymous):

lets say you have put an amount "P" in bank. question is asking to find how long will it take to double (2P) 2P = P * e^(.038*t) 2 = e^(.038*t) ln 2 = .038*t t = (ln 2)/.038

OpenStudy (anonymous):

Ok